# Quantum algorithms for calculating determinant and inverse of matrix and solving linear algebraic systems

### v3

@@ Abstract
We propose quantum algorithms for calculating the determinant and inverse of an $(N-1)\times(N-1)$ matrix (depth is $O(N^{2}\log N)$) which is a simple modification of the algorithm for calculating the determinant of an $N\times N$ matrix (depth is $O(N\log^{2}N)$. The basic idea is to encode each row of the matrix into a pure state of some quantum system. In addition, we use the representation of the elements of the inverse matrix in terms of algebraic complements. Combination of this algorithm with the quantum algorithm for matrix multiplication, proposed in our previous work, yields the algorithm for solving systems of linear algebraic equations (depth is $O(N\log^{2}N)$. Measurement of the ancilla state with output 1 (probability is $\sim 2^{-O(N\log N)}$) removes the garbage acquired during calculation. Appropriate circuits for all three algorithm are presented and have the same estimation $O(N\log N)$ for the size.
Keywords: quantum algorithm, quantum circuit, determinant, inverse matrix, linear system, depth of algorithm, quantum measurement
@@ I Introduction
Quantum algorithms that exploit principles of quantum physics can be run on quantum computers and promise to offer considerable advantages over classical counterparts. Algorithms for quantum calculations are an increasingly prominent field in quantum information research. Notable examples include Shor’s factoring algorithm Sho1 ; Sho2 , Deutch parallelism Deu , Grover’s search algorithm Gro , Quantum Fourier transform QFT1 ; QFT2 ; QFT3 and Quantum phase estimation Wang ; QFT3 . Further advancements include the HHL algorithm for solving systems of linear equations HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 , in which the crucial point is the conditional rotation of the state of the one-qubit ancilla. This operation, where inversion of the eigenvalue requires involving classical computing, is placed between the blocks of the direct and inverse phase estimation algorithms. HHL algorithm exponentiates a Hermitian matrix of the considered linear system, which can be performed via the Trotterization formula T ; AT . There is extensive literature devoted to improvement of this algorithm, especially its scale and complexity BHMT ; CKS ; GSLW ; LC ; LYC ; AL ; JRS ; SSO ; AlL ; TAWL . However, the principal problem of eigenvalue inversion is not resolved via quantum methods yet. For this purpose, for instance, the quantum state tomography can be involved to obtain the eigenvalues of the exponent of Hermitian operator (the matrix of the linear system) T . Nevertheless, the HHL algorithm combining the phase estimation and classical inversion of eigenvalues is very popular in realization of various algorithms of quantum computation and it is widely applicable in quantum machine learning A ; HHL2 ; RML . In particular, it is applied in the least-square linear-regression algorithms WBL ; SSP ; Wang2 working with large data sets, in quantum algorithms for singular-value decomposition GSLW ; G , in algorithms for solving linear B and nonlinear LKKLTC differential equations.
We also have to refer to algorithms for matrix manipulations based on the Trotterization method and the Baker-Champbell-Housdorff formula for exponentiating matrices ZhaoL1 . In GSLW , algebraic manipulations with matrices are implemented using unitary transformations that include the matrices as specific blocks (block-encoding model). A similar block-encoding approach was used in the algorithm of DFZ_2020 to embed the inverse of the matrix of an algebraic system into the unitary transformation. In LWWZ , an algorithm for matrix multiplication using binary encoding of matrix elements in the states of quantum systems followed by binary multiplication using Toffoli gates was proposed. Matrix multiplication over rings, including the multiplication of Boolean matrices was studied in KN .
@@
Later, in ZQKW_arxiv2023 , we proposed the implementation of unitary transformations corresponding to vector inner product, matrix sum, and product using multi-qubit conditional operators (reducible to Toffoli gates) and Hadamard transformations, including the ancilla measurement for garbage removal. The depth or complexity of these algorithms increases logarithmically with the matrix dimension, or even remains independent of it (in the case of matrix sum), demonstrating the advantage of these algorithms compared to their classical counterparts.
In this paper, we propose the algorithm for matrix inversion with its application to solving systems of linear equations. The principal part of this algorithm is the algorithm for computing the determinant of a square matrix. It is built on the ideas of Refs. QZKW_arxive2022 ; ZQKW_arxiv2023 :
    Encode matrix elements into the states of a certain quantum subsystems through probability amplitudes in such a way that all necessary multiplications of matrix elements are performed automatically and appear in the probability amplitudes of the initial state of the whole system;
@@
We follow the above steps to propose an algorithm for calculating the determinant of a square matrix whose $N$ rows are encoded into the pure states of $N$ separate quantum subsystems. Notice that creating such initial state is a special problem by itself, but we do not handle this problem here. The state of the whole system results in all possible products of $N$ elements from different rows. One only needs to select the proper products to calculate the determinant. We propose such an algorithm and construct the quantum circuit realizing it. While developing the algorithm, our focus is on the sizes and depths of its specific blocks. Essentially, our algorithm is a quantum realization of the cofactor expansion classical algorithm, whose depth exponentially increases with matrix dimension. However, the resulting depth of our algorithm is $O(N\log^{2}N)$ (or $O(N^{2}\log N)$ in the worst case) with the size $O(N\log N)$. The final step of this algorithm is the one-qubit ancilla measurement, which removes all the garbage. The probability of obtaining the required result of the ancilla measurement (which is 1) exponentially decreases with the matrix dimension $N$, which requires exponentially increasing number of evaluations of the algorithm. This is a central problem that we plan to overcome in further work.
@@
It is remarkable that the minor modification of the algorithm for calculating the determinant yields an alternative quantum algorithm for matrix inversion using its definition in terms of algebraic complements. We again use the row-wise encoding of the matrix elements into the state of a quantum system so that all necessary products of matrix elements appear automatically in the superposition state, thus one only needs to properly sort these products. The depth of this algorithm is $O(N^{2}\log N)$ (or $O(N^{2}\log^{2}N)$ in the worst case) and the size is $O(N\log N)$. Based on such inversion algorithm we develop an alternative algorithm for solving systems of linear algebraic equations. Thus, we do not use Trotterization methods and classical subroutine for inversion of eigenvalues. Therefore the proposed algorithm is completely quantum and it is an alternative to the inversion algorithm considered in HHL ; T ; TAWL . The estimations for the depth and size of this algorithm remain the same.
In relation with the question of algorithm complexity, we shall mention that the special methods of classical manipulations with matrices allow to reduce the depth of the algorithms for determinant calculation to $O(\log^{2}N)$ Ber ; BAEM with the size $O(N^{3.496})$. In our case we use the Leibniz formula as a definition of the matrix determinant and definition of the inverse matrix in terms of the algebraic complements. Our quantum algorithms based on these definitions obviously reveal advantages in comparison with classical realizations of these definitions. It is likely that using specific classical algorithms for matrix manipulations and developing their quantum analogues lead to further improvement in the depth and size of quantum algorithms calculating the determinant and inverse matrix. This issue is not considered in our paper. The main novelty is that we are promoting the method for representation of the matrix elements as probability amplitudes of pure states and reveal profits out of such representation in application to the standard algorithms for matrix manipulations.
@@
The paper is organized as follows. The algorithm for calculating the matrix determinant is proposed in Sec.II. This algorithm is taken as a basis for the algorithm calculating the inverse matrix presented in Sec. III. Based on that results, the algorithm for solving systems of linear algebraic equations is proposed in Sec. IV. The main results are discussed in Sec.V. Appendix, Sec.VI, contains the modified algorithms for matrix multiplication.
@@ II Determinant
we encode each row of $N$ elements into the pure state $|\Psi_{j}\rangle$ of the $n$-qubit subsystem $S_{j}$,
@@
Representing rows of the matrix $\tilde{A}$ in the form of pure states (7) presumes $N$ obvious constraints for the matrix elements $a_{ij}$ predicted by the normalization of pure states:
@@
These constraints can be smoothened by incorporating additional terms with different probability amplitudes into the states $|\Psi_{i}\rangle$, but this issue will be addressed in future study. The whole system is the composition of the subsystems, $S=\bigcup_{j}S_{j}$, and the pure state of the whole system reads
@@
$=$ $\sum_{k_{0},\dots,k_{N-1}=0}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}},$
@@
$|\Phi_{0}^{(D)}\rangle$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle,$ (10)
where $|g_{1}\rangle$ collects all extra terms which form garbage to be removed later.
We have to arrange the appropriate signs for each term of the first part in the right-hand side of (10). This can be achieved by setting the subsystem $S_{j}$ (in each term) into the state $|j\rangle$ using the SWAP operators together with sign-switching. This pair of operators performs permutation properly altering the sign of the term.
Thus, our purpose is to exchange the states of subsystems $S_{j}$ ($j=0,\dots,N-1$) so that the state of the subsystem $S_{k}$ becomes $|k\rangle_{S_{k}}$ in each term of the first part of superposition state (10). For that we introduce the projectors $P_{S_{j}}^{(k)}=|k\rangle_{S_{j}}\;{}_{S_{j}}\langle k|$, ($k=0,\dots N-2,$ $j=k+1,\dots N-1$) and the $\tilde{n}_{k}$-qubit ancillae $A_{k}$, ($\tilde{n}_{k}=\lceil\log(N-k)\rceil$, $k=0,\dots N-2$). The ancilla $A_{k}$ serves to save the index $j$ of the subsystem $S_{j}$ whose state is $|k\rangle$ with $j>k$. Therefore, the number of qubits $\tilde{n}_{k}$ in $A_{k}$ is less then $n=\log N$ unless $k=0$: $\tilde{n}_{k}\leq n$. For instance, in Fig. 2 for the case $N=4$, there are three ancillae $A_{0},A_{1},A_{2}$. Ancilla $A_{2}$ has to encode either only one state $|3\rangle$ or nothing, therefore one qubit is enough: we can encode $|3\rangle$ by the state $|1\rangle$ of ancilla $A_{2}$. Next, $A_{1}$ has to encode two states $|2\rangle$ and $|3\rangle$ or nothing. Two qubits are enough for that: $|2\rangle\rightarrow|1\rangle,|3\rangle\rightarrow|2\rangle$. Similarly, $A_{0}$ requires 2 qubits as well. Using the projectors $P_{S_{j}}^{(k)}$ and ancillae $A_{k}$ we can prepare the operators $V_{S_{j}A_{k}}$, which relates the index $j$ of $S_{j}$ to the appropriate state $|\tilde{j}\rangle_{A_{k}}$ of the ancilla $A_{k}$:
@@
where $\sigma^{(x)}_{kj}$ includes the set of $\sigma^{(x)}$-operators applied to the ground state of the ancilla $A_{k}$ to form the state $|\tilde{j}\rangle_{A_{k}}$; $Z_{k}(j)$ changes the sign of the ancilla state $|\tilde{j}\rangle_{A_{k}}$ (it counts permutations and can be represented by either the $\sigma^{(z)}$-operator applied to one of the excited qubits, or by the $\sigma^{(x)}\sigma^{(z)}\sigma^{(x)}$-operator applied to one of the ground-state qubits (only to one!) of $A_{k}$). To arrange the exchange of states between $S_{j}$ and $S_{k}$, we introduce the projectors $P_{A_{k}}^{(j)}=|\tilde{j}\rangle_{A_{k}}\;{}_{A_{k}}\langle\tilde{j}|,$ $k=0,\dots N-2,$ $j=k+1,\dots N-1,$ $\tilde{j}=0,\dots N-k-2,$ and construct the operators $U_{S_{k}S_{j}}$:
@@
Figure 1: $V$-and $U$-operators: (a) the general case; (b) the $4\times 4$-matrix case, then $\tilde{n}_{0}=2$, $\tilde{n}_{1}=\tilde{n}_{2}=1$, $\tilde{n}_{k}<n$, therefore, $A_{1}$ is a one-qubit ancilla.
The depth of each operator $V_{S_{j}A_{k}}$ is $O(n)$ (due to the projectors $P_{S_{j}}^{(k)}$ with $n$ control qubits and we assume that all $\sigma^{(x)}$-operators of $\sigma^{(x)}_{kj}$ are applied in parallel to different qubits of $a_{k}$), while the depth of $U_{S_{k}S_{j}}$ is the of depth of the control-SWAP-operator which includes $n$ swaps each controlled by $n$ qubits, i.e. the total depth is $n\cdot n=n^{2}$ (or $O(n^{2})$).
We prepare the operators
$W_{k}=\prod_{j=k+1}^{N-1}U_{S_{k}S_{j}}V_{S_{j}A_{k}}.$ (13)
Applying $W_{k}$ to the state $|\Phi_{0}\rangle|0\rangle_{A}$ we put $S_{k}$ into the state $|k\rangle_{S_{k}}$ in each term of the first part of (10). Although $W_{k}$ has product of $N-k-1$ pairs of $U$- and $V$-operators, only one pair of such operators is applied to each particular term in $|\Phi_{0}\rangle$ which is due to the structure of the projector operators $P^{(k)}_{S_{j}}$. We assume that this means the $UV$-products can be applied in parallel and thus the depth of $W_{k}$ is $O(n)+O(n^{2})=O(n^{2})$. Otherwise we have to multiply the obtained depth by $O(N)$. The operator $\prod_{k=0}^{N-2}W_{k}$ provides the needed sign of the probability amplitudes which will compose the expression for the determinant. To collect all of them in a single probability amplitude we apply the Hadamard transformation to each qubit of ancillae $A_{k}$, $k=0,\dots,N-2$, and denote this transformations $H_{A}$. There are $\tilde{N}$ Hadamard operators in $H_{A}$,
@@
$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{2}\rangle_{SA},$
@@
$\det(A)$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}\epsilon_{k_{0}\dots k_{N-1}}a_{k_{0}0}\dots a_{k_{N-1}(N-1)},$ (17)
and $\epsilon_{k_{0}\dots k_{N-1}}$ is the permutation symbol. The depth of $W^{(1)}$ is $O(Nn^{2})$. Next, we can set states of all $S_{j}$ in the first term of $|\Phi_{1}\rangle$ into the $|0\rangle_{S_{j}}$ states. To this end we apply the $\sigma^{x}$-operators to the excited qubits of $S$, i.e. prepare the operator $X_{S}$,
@@
with $X_{S_{j}}(j)$ being the set of $\sigma^{(x)}$ operators acting on the excited qubits of $S_{j}$ in the state $|j\rangle_{S_{j}}$. The depth of $X_{S}$ is $O(1)$:
@@
$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}+|g_{3}\rangle_{SA}.$
We notice that the operators $H_{A}$ in $W^{(1)}_{SA}$ and operator $X_{S}$ can be applied simultaneously because they effect on different qubits. We separate them to underline the distinction in their functions.
Next, we remove the garbage. For that we introduce the projector $P_{SA}=|0\rangle_{S}|0\rangle_{A}\;{}_{A}\langle 0|_{S}\langle 0|$, include the one-qubit ancilla $B$ in the ground state and construct the operator $W^{(2)}_{SAB}$,
@@
$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g_{3}\rangle_{SA}|0\rangle_{B}.$
Now we measure the ancilla $B$ with the output $|1\rangle_{B}$. The probability of this result of measurement is $2^{-\tilde{N}}|{\det(A)}|^{2}$. After the measurement we get the following state:
$|\Phi_{4}^{(D)}\rangle=|\Psi^{(D)}_{out}\rangle|0\rangle_{A},\;\;|\Psi^{(D)}_{out}\rangle=\arg(\det(A))\;|0\rangle_{S},$ (22)
The particular circuit for calculating the $4\times 4$ determinant using the proposed algorithm is given in Fig. 2. The circuit for the determinant of $N\times N$ matrix includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the subsystem $A$ and one qubit of the ancilla $B$, therefore the size of this circuit is $O(N\log N)$.
Figure 2: Quantum circuit illustrating computation of the determinant of a $4\times 4$ matrix. We omit superscript $(D)$ at $|\Phi_{k}\rangle$, $k=1,\dots,4$, and at $|\Psi_{out}\rangle$ for brevity. Operators $H_{A}$ and $X_{S}$ can be applied in parallel since they effect on different qubits.

### v4

# Purely quantum algorithms for calculating determinant and inverse of matrix and solving linear algebraic systems

@@ Abstract
We propose quantum algorithms that are fundamentally quantum in nature for the computation of the determinant and inverse of an $(N-1)\times(N-1)$ matrix, with a computational depth of $O(N^{2}\log N)$. This approach is a straightforward modification of the existing algorithm for determining the determinant of an $N\times N$ matrix, which operates with a depth of $O(N\log^{2}N)$. The core concept involves encoding each row of the matrix into a pure state of a quantum system. In addition, we use the representation of the elements of the inverse matrix in terms of algebraic complements. This algorithm, in conjunction with the matrix multiplication algorithm discussed in the appendix, facilitates solving systems of linear algebraic equations with a depth of $O(N\log^{2}N)$. The measurement of the ancilla state yielding an output of 1 occurs with a probability of approximately $2^{-O(N\log N)}$, and eliminates the extraneous data generated during the computation. We provide suitable circuit designs for all three algorithms, each estimated to require $O(N\log N)$ in terms of space, specifically the number of qubits utilized in the circuit.
Keywords: quantum algorithm, quantum circuit, determinant and inverse of matrix, linear algebraic systems, depth of algorithm, quantum measurement
@@ I Introduction
Quantum algorithms exploiting principles of quantum physics can be run on quantum computers and promise to offer considerable advantages over classical counterparts. Algorithms for quantum calculations are an increasingly prominent field in quantum information research. Notable examples include Shor’s factoring algorithm Sho1 ; Sho2 , Deutsch parallelism Deu , Grover’s search algorithm Gro , Quantum Fourier transform QFT1 ; QFT2 ; QFT3 and Quantum phase estimation Wang ; QFT3 . Further advancements include the HHL algorithm for solving systems of linear equations HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 , in which the crucial point is the conditional rotation of the state of the one-qubit ancilla. This operation, where inversion of the eigenvalue requires involving classical computing, is placed between the blocks of the direct and inverse phase estimation algorithms. HHL algorithm exponentiates a Hermitian matrix of the considered linear system, which can be performed via the Trotterization formula T ; AT . There is extensive literature devoted to improvement of this algorithm, especially its scaling and complexity BHMT ; CKS ; GSLW ; LC ; LYC ; AL ; JRS ; SSO ; AlL ; TAWL . However, the principal problem of eigenvalue inversion is not resolved via quantum methods yet. For this purpose, for instance, the quantum state tomography can be used to obtain the eigenvalues of the exponent of Hermitian operator (the matrix of the linear system) T . Nevertheless, the HHL algorithm combining the phase estimation and classical inversion of eigenvalues is very popular in implementation of various algorithms of quantum computation and it is widely applicable in quantum machine learning A ; HHL2 ; RML . In particular, it is applied in the least-square linear-regression algorithms WBL ; SSP ; Wang2 working with large data sets, in quantum algorithms for singular-value decomposition GSLW ; G , in algorithms for solving linear B and nonlinear LKKLTC differential equations.
We also refer to algorithms for matrix manipulations based on the Trotterization method and the Baker-Champbell-Hausdorff formula for exponentiating matrices ZhaoL1 . In GSLW , algebraic manipulations with matrices are implemented using unitary transformations that include the matrices as specific blocks (block-encoding model). A similar block-encoding approach was used in the algorithm of DFZ_2020 to embed the inverse of the matrix of an algebraic system into the unitary transformation. In LWWZ , an algorithm for matrix multiplication using binary encoding of matrix elements in the states of quantum systems followed by binary multiplication using Toffoli gates was proposed. Matrix multiplication over rings, including the multiplication of Boolean matrices was studied in KN .
@@
Later, in ZQKW_arxiv2023 , we proposed the implementation of unitary transformations corresponding to vector inner product, matrix sum, and product using multi-qubit conditional operators (reducible to Toffoli gates) and Hadamard transformations, including the ancilla measurement for garbage removal. The depth of these algorithms increases logarithmically with the matrix dimension, or even remains independent of it (in the case of matrix sum), demonstrating the advantage of these algorithms compared to their classical counterparts.
In this paper, we propose quantum algorithms for determinant and inversion of matrix with their application to solving systems of linear equations. The principal part of these algorithms is the algorithm for computing the determinant of a square matrix. It is built on the ideas of Refs. QZKW_arxive2022 ; ZQKW_arxiv2023 :
    Encode matrix elements into the states of certain quantum subsystems through probability amplitudes in such a way that all necessary multiplications of matrix elements are performed automatically and appear in the probability amplitudes of the initial state of the whole system;
@@
We follow the above steps to propose an algorithm for calculating the determinant of a square matrix whose $N$ rows are encoded into the pure states of $N$ separate quantum subsystems. Notice that creating such initial states is a special problem in itself, but we do not discuss this problem here. The state of the whole system results in all possible products of $N$ elements from different rows. One only needs to select the proper products to calculate the determinant. We propose such an algorithm and construct the quantum circuit realizing it. While developing the algorithm, our focus is on the space and depths of its specific blocks. Essentially, our algorithm is a quantum realization of the cofactor expansion of the classical algorithm, whose depth exponentially increases with matrix dimension. However, the resulting depth of our algorithm is $O(N\log^{2}N)$ (or $O(N^{2}\log N)$ in the worst case) with space of $O(N\log N)$ qubits. The final step of this algorithm is the one-qubit ancilla measurement which removes all the garbage. The probability of obtaining the required result of the ancilla measurement (which is 1) diminishes exponentially as the matrix dimension $N$ increases, necessitating an exponentially larger number of algorithm evaluations. This is a central problem that we plan to investigate in future work.
@@
It is remarkable that the minor modification of the algorithm for calculating the determinant yields an alternative quantum algorithm for matrix inversion using its definition in terms of algebraic complements. We again use the row-wise encoding of the matrix elements into the state of a quantum system so that all necessary products of matrix elements appear automatically in the superposition state. One thus only needs to properly sort these products. The depth of this algorithm is $O(N^{2}\log N)$ (or $O(N^{2}\log^{2}N)$ in the worst case) and the space is $O(N\log N)$ qubits. Based on such inversion algorithm we develop an alternative algorithm for solving systems of linear algebraic equations. Thus, we do not use Trotterization methods and classical subroutine for inversion of eigenvalues. Therefore, the proposed algorithm is purely quantum in nature and it gives an alternative avenue to the inversion algorithm considered in HHL ; T ; TAWL . The estimations for the depth and space of this algorithm remain the same.
Regarding the question of algorithmic complexity, we remark that the specialized classical techniques of matrix manipulations facilitate a reduction in the depth of the algorithms for determinant calculation to $O(\log^{2}N)$ Ber ; BAEM with the size $O(N^{2.496})$. In our study, we employ the Leibniz formula as a definition for the matrix determinant and define the inverse matrix in terms of the algebraic complements. Our quantum algorithms based on these definitions demonstrate clear advantages over their classical counterparts. It is plausible that the adaptation of specific classical algorithms for matrix operations into their quantum equivalents could yield further enhancements in the efficiency and complexity of quantum algorithms designed for calculating determinants and inverse matrices. However, this particular aspect is not addressed in the present paper. The primary contribution of our work lies in promoting the representation of matrix elements as probability amplitudes of pure states, and we elucidate the benefits of this representation in relation to standard algorithms for matrix manipulations.
@@
The paper is organized as follows. The algorithm for calculating the matrix determinant is proposed in Sec.II. This algorithm is taken as a basis for the algorithm calculating the inverse matrix presented in Sec.III. Based on those results, the algorithm for solving systems of linear algebraic equations is proposed in Sec.IV. The main results are discussed in Sec.V. Appendix, Sec.VI, contains the modified algorithms for matrix multiplication.
@@ II Determinant
we encode each row consisting of $N$ elements into the pure state $|\Psi_{j}\rangle$ of the $n$-qubit subsystem $S_{j}$,
@@
Representing rows of the matrix $\tilde{A}$ in the form of pure states (7) creates $N$ obvious constraints for the matrix elements $a_{ij}$ because of the normalization of pure states:
@@
These constraints can be circumvented by incorporating additional terms with different probability amplitudes into the states $|\Psi_{i}\rangle$, but this issue will be addressed in future study. The whole system is the union of the subsystems, $S=\bigcup_{j}S_{j}$, and the pure state of the whole system reads
@@
$=$ $\sum_{k_{0},\dots,k_{N-1}=0}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}$
$\times$ $|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}},$
@@
$|\Phi_{0}\rangle$ $=$ $\sum_{{k_{0},\dots,k_{N-1}=0}\atop{k_{0}\neq\dots\neq k_{N-1}}}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}$ (10)
$\times$ $|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle_{S},$
where $|g_{1}\rangle$ collects all extra terms which form the garbage to be removed later. We have to arrange the appropriate signs for each term of the first part in the right-hand side of (10). This can be achieved by setting the subsystem $S_{j}$ (in each term) into the state $|j\rangle$ using the SWAP operators together with sign-switching. This pair of operators performs permutation and changes the sign of the term.
Thus, our purpose is to exchange the states of subsystems $S_{j}$ ($j=0,\dots,N-1$) so that the state of the subsystem $S_{k}$ becomes $|k\rangle_{S_{k}}$ in each term of the first part of superposition state (10). For that we introduce the projectors $P_{S_{j}}^{(k)}=|k\rangle_{S_{j}}\;{}_{S_{j}}\langle k|$, ($k=0,\dots N-2,$ $j=k+1,\dots N-1$) and the $\tilde{n}_{k}$-qubit ancillae $A_{k}$, ($\tilde{n}_{k}=\lceil\log(N-k)\rceil$, $k=0,\dots N-2$). The ancilla $A_{k}$ serves to save the index $j$ of the subsystem $S_{j}$ whose state is $|k\rangle$ with $j>k$. Therefore, the number of qubits $\tilde{n}_{k}$ in $A_{k}$ is less than $n=\log N$ unless $k=0$: $\tilde{n}_{k}\leq n$. For instance, in Fig. 2 for the case $N=4$, there are three ancillae $A_{0},A_{1},A_{2}$. Ancilla $A_{2}$ has to encode either only one state $|3\rangle$ or nothing, therefore one qubit is enough: we can encode $|3\rangle$ by the state $|1\rangle$ of ancilla $A_{2}$. Next, $A_{1}$ has to encode two states $|2\rangle$ and $|3\rangle$ or nothing. Two qubits are enough for that: $|2\rangle\rightarrow|1\rangle,|3\rangle\rightarrow|2\rangle$. Similarly, $A_{0}$ requires 2 qubits as well. Using the projectors $P_{S_{j}}^{(k)}$ and ancillae $A_{k}$ we can prepare the operators $V_{S_{j}A_{k}}$, which relates the index $j$ of $S_{j}$ to the appropriate state $|\tilde{j}\rangle_{A_{k}}$ of the ancilla $A_{k}$:
@@
where $\sigma^{(x)}_{kj}$ includes the set of $\sigma^{(x)}$-operators which change the ground state of the ancilla $A_{k}$ to the state $|\tilde{j}\rangle_{A_{k}}$; $Z_{k}(j)$ changes the sign of the ancilla state $|\tilde{j}\rangle_{A_{k}}$ (it counts swaps and can be represented by either the $\sigma^{(z)}$-operator applied to one of the excited qubits, or by the $\sigma^{(x)}\sigma^{(z)}\sigma^{(x)}$-operator applied to one of the ground-state qubits (only to one!) of $A_{k}$). To arrange the exchange of states between $S_{j}$ and $S_{k}$, we introduce the projectors $P_{A_{k}}^{(j)}=|\tilde{j}\rangle_{A_{k}}\;{}_{A_{k}}\langle\tilde{j}|,$ $k=0,\dots N-2,$ $j=k+1,\dots N-1,$ $\tilde{j}=0,\dots N-k-2$, $\tilde{j}=j-k-1,$ and construct the operators $U_{S_{k}S_{j}}$:
@@
Figure 1: $V$ and $U$ operators for (a) the general case and (b) the $4\times 4$-matrix case. In the latter case, $\tilde{n}_{0}=\tilde{n}_{1}=2$, $\tilde{n}_{2}=1$, $\tilde{n}_{k}<n$. Hence, $A_{1}$ is a two-qubit ancilla.
The depth of each operator $V_{S_{j}A_{k}}$ is $O(n)$ (due to the projectors $P_{S_{j}}^{(k)}$ with $n$ control qubits and we assume that all $\sigma^{(x)}$-operators of $\sigma^{(x)}_{kj}$ are applied in parallel to different qubits of $A_{k}$), while the depth of $U_{S_{k}S_{j}}$ is the depth of the control-SWAP-operator which includes $n$ swaps each controlled by $n$ qubits, i.e., the total depth is $nO(n)=O(n^{2})$. Theoretically, it can be implemented with less operations but this would not change the asymptotic complexity of the circuit.
Applying the operators
$W_{k}=\prod_{j=k+1}^{N-1}U_{S_{k}S_{j}}V_{S_{j}A_{k}},$ (13)
to the state $|\Phi_{0}\rangle|0\rangle_{A}$, we put the subsystem $S_{k}$ into the state $|k\rangle_{S_{k}}$ in each term of the first part of (10). Although $W_{k}$ has product of $N-k-1$ pairs of $U$- and $V$-operators, only one pair of such operators is applied to each particular term in $|\Phi_{0}\rangle$ which is due to the structure of the projector operators $P^{(k)}_{S_{j}}$. We assume that this means the $UV$-products can be applied in parallel and thus the depth of $W_{k}$ is $O(n)+O(n^{2})=O(n^{2})$. Otherwise we have to multiply the obtained depth by $O(N)$. The operator $\prod_{k=0}^{N-2}W_{k}$ provides the needed sign of the probability amplitudes which will compose the expression for the determinant. To collect all of them in a single probability amplitude we apply the Hadamard transformation to each qubit of ancillae $A_{k}$, $k=0,\dots,N-2$, and denote this transformations $H_{A}$. There are $\tilde{N}$ Hadamard operators in $H_{A}$,
@@
$=\frac{\det(\tilde{A})}{2^{\tilde{N}/2}}\;|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{2}\rangle_{SA},$
@@
$\det(\tilde{A})$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}\epsilon_{k_{0}\dots k_{N-1}}a_{k_{0}0}\dots a_{k_{N-1}(N-1)},$ (17)
and $\epsilon_{k_{0}\dots k_{N-1}}$ is the permutation symbol. The depth of $W^{(1)}$ is $O(Nn^{2})$. Next, we can set states of all $S_{j}$ in the first term of $|\Phi_{1}^{(D)}\rangle$ into the $|0\rangle_{S_{j}}$ states. To this end we apply the $\sigma^{x}$-operators to the excited qubits of $S$, i.e. prepare the operator $X_{S}$.
@@
with $X_{S_{j}}(j)$ being the set of $\sigma^{(x)}$ operators acting on the excited qubits of $S_{j}$ in the state $|j\rangle_{S_{j}}$. The depth of $X_{S}$ is $O(1)$. Applying $X_{S}$ to $|\Phi_{1}^{(D)}\rangle$ we obtain:
@@
$=\frac{\det(\tilde{A})}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}+|g_{3}\rangle_{SA}.$
We note that the operators $H_{A}$ in $W^{(1)}_{SA}$ and operator $X_{S}$ can be applied simultaneously because they affect different qubits. We separate them to underline the distinction in their functions. Next, we remove the garbage. For that we introduce the projector $P_{SA}=|0\rangle_{S}|0\rangle_{A}\;{}_{A}\langle 0|_{S}\langle 0|$, include the one-qubit ancilla $B$ in the ground state and construct the operator $W^{(2)}_{SAB}$,
@@
$=\frac{\det(\tilde{A})}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g_{3}\rangle_{SA}|0\rangle_{B}.$
Now we measure the ancilla $B$ with the output $|1\rangle_{B}$. The probability of this result of measurement is $2^{-\tilde{N}}|{\det(\tilde{A})}|^{2}$. After the measurement we get the following state:
$|\Phi_{4}^{(D)}\rangle=|\Psi^{(D)}_{out}\rangle|0\rangle_{A},\;\;|\Psi^{(D)}_{out}\rangle=\arg(\det(\tilde{A}))\;|0\rangle_{S},$ (22)
The particular circuit for calculating the $4\times 4$ determinant using the proposed algorithm is given in Fig. 2. The circuit for the determinant of $N\times N$ matrix includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the subsystem $A$ and one qubit of the ancilla $B$, therefore the space of this circuit is $O(N\log N)$ qubits.
Figure 2: Quantum circuit illustrating computation of determinant of a $4\times 4$ matrix. We omit superscript $(D)$ in $|\Phi_{k}\rangle$ $(k=1,\dots,4)$ and $|\Psi_{out}\rangle$ for brevity. Operators $H_{A}$ and $X_{S}$ can be applied in parallel since they affect different qubits.